Hi

What are the products of this reaction:

B + C + LiNO3 + O2 ———->

Thanks.
1
Expert's answer
2018-10-08T07:48:08-0400

B + C + LiNO3 + O2 = >

4B + 3O2 = 2B2O3 (7000 С)

C + O2 = CO2 (600 -7000 С)


LiNO3 does not interact with other elements because the nitrogen in this salt has the highest degree of oxidation.